ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: Dr. Aboubakr Bayoumi, King Saud University College of Science, Mathematics Dept., P. O. Box 2455, Riyadh 11451 Saudi Arabia Quasi-Differentiability and Applications in Locally Bounded Spaces We introduce and study a new concept of differentiability, which we call “Quasi differentiability”, of mappings f between locally bounded spaces. This concept has wide applications in different branches of Mathematics such as Functional Analysis and Complex Analysis. As a consequence we obtained a new version of the Lagrange Mean-Value Theorem and Fundamental Theorem of Calculus in locally bounded spaces.